Tohban Report 2015-04-01

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 25 Mar 2015
End Date: 01 Apr 2015
Tohban: Milo Buitrago-Casas
Tohban email:
Next Tohban: Lindsay Erin Glesener
List all reports


Solar Activity

Solar activity was rather low during last days. There were not a single M or X class flare, but we sporadically saw several C-class flares. Currently there is only one active region on the disk close to the East limb that will go out of disk in one day and a half days. Solar activity is expected to be low during the following days.

How many GOES flares occurred?

 Flares above B, C, M, X class were      9    42     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      8    35     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     0     0     0

There were RHESSI flares/GOES flares 370 / 51 over the time range 24-Mar-15 31-Mar-15

Memory Management

The SSR was oscillating rawly between 10 and 35% all the week. Currently we are in an Active/Vigorous decimations but due to the low solar activity we recommend to the next Tohban to change it to Normal/Vigorous. Nevertheless it is recommended to keep a close eye on the SSR mainly because some thresholds of the detectors were changed during last days.

Spacecraft Status

Several changes were set on detector thresholds and high voltages last week. In general all detectors are stable but detector 9 was set at the maximum threshold allowed and at very low value of the high voltage. Also Detector 6 is showing a high value on the front slow count rates. All the other detectors are working well as far as the Tohban can tell.

A cryocooler experiment was carried out last Monday March 30. The current cryocooler power is at 75W (it was 71W until last week). We hope to discuss today

Data Gaps

There was no significant data gaps.


We increased the cryocooler power from around 71W to 75W last Monday to see if we could significantly drop down the Tips Temperatures. As a result during the following two days we observed a reduction on each temperature of around 0.6 degrees with the cryocooler constantly running at that relatively high power. Other parameters as the acceleration slightly raised up but kept stable during the time of the experiment. We need to discuss during the RHESSI meeting what to do with power. The low change in the tip temperature that we got does not justify the high power that we are using now. We strongly recommend go back to a lower power.

Detector issues

On March 25 tarting at 19:26 UTC we made a few changes to detector 6, 8 and 9. We decreased the HV on detectors 6 and 9 and we increased the rear fast thresholds on detectors 8 and 9.

Changes on High Voltages for detectors 6 and 9:

High Voltage [G6] Counts Voltage
Initial 213 4142 V
Final 193 3750 V

High Voltage [G9] Counts Voltage
Initial 186 3603 V
Final 136 2622 V

Both, detectors 6 and 9, remained segmented after the voltage changes

Changes of rear fast thresholds on detectors 8 and 9:

Threshold [G8] Rear fast
Initial 0x55
Final 0x65
Threshold [G9] Rear fast
Initial 0xD0
Final 0xFF

After this changes rear live-times raised up above 90%.

Only detectors 2 and 4 continue unsegmented.

And as it has been happening recently, there is still a light leak on detector 2.

Other notes

We suggest to the next Tohban to keep an eye on Tip temperatures, as well as to observe the evolution of detector 9 and 6.

Spacecraft Management

Decimation Active/Vigorous
Night time data (fronts) full nighttime events
Night time data (rears) full nighttime events
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Personal tools